OpenStudy (anonymous):

Find the slope of the line that contains the following points. R(-3, 5), S(3, -2)

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = ( -2- 5)/(3-(-3)) I'll let you finish

OpenStudy (anonymous):

undefined?

OpenStudy (anonymous):

7/6?

OpenStudy (anonymous):

thanks!!!

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( -2- 5)/(3-(-3)) m = ( -2- 5)/(3+3) m = -7/6

jimthompson5910 (jim_thompson5910):

yw